(1) Find and classify the extremes value of f(x) = 6x4 - 4x6 over the interval [ -2, 2].


(2) A retailer determines that the cost of ordering and storing units of a product can be modelled by C(x) = 3x + (2000/x), 0 ≤ x ≤ 200. Find the order size that will minimize ordering and storage cost. 


1
Expert's answer
2021-11-19T15:27:34-0500

2)

"C'(x) = 3x + (2000\/x)=3-2000\/x^2=0"

order size that will minimize ordering and storage cost:

"x=\sqrt{2000/3}=25.81" units


1)

"f(-2)=f(2) = 6x^4 - 4x^6=-160"


"f'(x)=24x^3-24x^5=0"

"x_1=0,x_2=1,x_3=-1"

"f(0)=0,f(1)=f(-1)=2"

so, there is absolute minimum at points (-2,-160) and (2,-160)

and absolute maximum at points (-1, 2) and (1, 2)

at point (0,0) derivative change sign from - to +, so it is local minimum